Breakdown Field (drawn from the text by Giancoli) Dry spo concentric with the sphere. In particular, use: Brew bre down and generate a spark if the electric feld exceseds about 3 x 10* N/C. low much charge could be packed onto a green pea (diameter 0.75 em) before the pea ntaneously discharges? It may be useful to take advantage that for uniform spherical charge distributions, the electric field outside the sphere is exactly that of a point charge O A. 5.1x 10-6 c ○ B, 3.7 × 10-7 C OC, 6.2 × 10-8 C a D. 4.7 × 10-9 c 0.00375) (3moetooo 375) 4x 109 Pea #8 Vanishing Electric Field The separation between a 1.00 μ0 point charge and an 8 00 /.c point charge is 2.0 meters. How far from the 1.00 does the electric field vanish? ○ A. 0.328 m ○ B. 0.522 m ° C. 0.853 m ○ D. 1.43m #9 Uniformly charged sphere A sphere has a total charge of 1.60 JC spread uniformly through its volume. If the sphere has a radius of 0.250 meters, what is the electric field at the spheres surface? 0 A. 0.00 N/C 0 B. 1.06 × 105 N/C C. 1.74 × 105 N/C D. 2.30 × 10s N/C - 2.3 x 105 #10 Uniformly charged Sphere interior In #9, what is the electric field strength in the interior of the sphere a distance of 0.125 meters from the center? A. 0.00 N/C 0 B. 1.15 x 105 N/C ° C. 176 × 105 N/C D. 2.30 × 105 N/C
Answer #1

8.

r = distance between the two charges = 2 m

d = distance of the point from 1 uC

for the electric field to vanish at the point

E1 = E2

k Q1/d2 = k Q2/(r - d)2

1/d2 = 8/(2 - d)2

d = 0.522 m